A jet airliner flies at a constant speed, covering 401 miles in 1.8 hours. What is the speed of the plane in miles per hour?


1
Expert's answer
2021-04-13T06:42:01-0400
"v=\dfrac{s}{t}=\dfrac{401\ mi}{1.8\ h}=223\ \dfrac{mi}{h}."